OpenStudy (msmr):

Find domain and range of f(x)=sqrt x+3?

OpenStudy (msmr):

just to clear it up: \[f(x) = \sqrt{x+3}\]

OpenStudy (anonymous):

\(x+3 \ge 0 \implies x \ge -3 \implies x\in [-3,\infty)\)

OpenStudy (msmr):

that's the range, right?

OpenStudy (msmr):

what about 1/(x+3)^2?

OpenStudy (anonymous):

That's domain.

OpenStudy (anonymous):

Why don't you post it as a new question?